# Metaverse Identity: Core Principles and Critical Challenges

### v1

@@ 3.1 Equivalence and Alignment
In recent years, the incidence of sexual assaults within the metaverse has become as a pressing concern, presenting complex social and legal dilemmas (Schulenberg et al.,, 2023; Lee et al., 2023b, ; Bellini,, 2024). For instance, in December 2021, a female Metaverse user reported that three male avatars had ’touched her[her avatar] inappropriately’ (Clayton,, 2022). Another disturbing incident occurred in April 2022, when a female user from Japan shared via Twitter her experience of being attacked and virtually raped in VRChat while she was asleep (Anonymous,, 2022). Additionally, the Daily Mail in the UK highlighted a particularly alarming case in January with its headline "British police probe Virtual Rape in metaverse: Young girl’s digital persona ’is sexually attacked by gang of adult men in immersive video game’ - sparking the first investigation of its kind and questions about extent current laws apply in online world" (Camber Rebecca,, 2024). The incident involved a British girl under the age of 16 whose avatar was raped by multiple avatars controlled by strangers during a virtual reality game, leading to the first global investigation of such behaviors and raising significant legal inquiries regarding the applicability of existing laws in the metaverse.
@@
At first glance, these cases may seem to involve merely disputes over unethical conduct and criminal actions, along with attendant legal applicability issues. However, a deeper analysis reveals that they fundamentally reflect profound challenges associated with identity recognition in the metaverse (Cheong,, 2022; Kasiyanto et al.,, 2022; Wu and Zhang,, 2023; Lee et al., 2023b, ). The reason these incidents of sexual assault attract world-spread social attention and provoke strong reaction lies in the deep psychological identification that individuals have with their avatars (Lee et al., 2023b, ). If avatars were merely considered gaming tools or entertainment elements, violations against them would not significantly impact users in reality. Nevertheless, the victims’ report of discomfort and psychological trauma suggest a profound association between their avatar and their identities. This strong identification means that harm inflicted on an avatar is perceived as harm to the self, eliciting genuine emotional and psychological responses (Lee et al., 2023b, ). The legal disputes triggered by these cases also underscore the unique nature of metaverse identities in both legal and social contexts. While violations of personal rights in the physical world inevitably entail legal repercussions, in the metaverse, the identity attributes of avatars remain defined, leading to ambiguities in rights protection (Lee et al., 2023b, ; Bellini,, 2024). This ambiguity necessitates a redefinition of the identities represented by avatars. Such incidents of ’sexual assault’ not only reveal the high degree of integration between avatars and users’ self-presence-a psychological state in which people equate their virtual selves with their actual selves (Lee et al., 2023b, )- but also demonstrate the significant influence of metaverse identities on real-life identity, highlighting the close connection between the metaverse and real life.
@@
It is critical to acknowledge that the identity dilemmas triggered by these ’sexual assault’ cases represent just one of the numerous challenges posed by metaverse entities in the metaverse. OpenAI’s recent release of ChatGPT-4o [^2] https://openai.com/index/hello-gpt-4o/, an AI chatbot capable of human-like interaction, introduces a novel and complex dimension of challenges for identity recognition in the metaverse (Hsu,, 2024). In this context, we propose the core principle - Equivalence and Alignment for conceptualizing and governing Metaverse Identity. Alignment refers to the consistency between metaverse identities and real-world identities concerning behavioral norms, social guidelines, and ethical standards. Equivalence emphasizes the parity between the two in aspects such as behavioral patterns, emotional experiences, cognitive perceptions, and economic interests (Lee et al., 2023b, ). Essentially, Alignment mandates that metaverse identities adheres to the same standards as real-world identities, while Equivalence underscores that the impact of virtual identities on users is as significant as that of real-world identities. We aim to simplify the complex attribute of Metaverse Identity through this principle, providing a clear and straightforward framework for developing metaverse identities. The principle also underscores the unique status of Metaverse Identity in contemporary legal and social contexts in metaverse era compared to in internet era, indicating the necessity of refining relevant regulations in the physical world. Furthermore, we intend to offer principled recommendations for policymakers, technology developers, and platform operators. While striving for deeper immersion and realism in the metaverse, it is imperative to prioritize the recognition and protection of user identities, thereby guiding necessary regulation and ethical oversight at the intersection of technology and ethics.
@@ 3.2 Fusion and Expansiveness
Hatada et al., (2024) recently published their study "Avatar Robot Cafe," which explores how individuals with disabilities employ robots, avatars and a hybrid cyber-physical environment to redefine their identities. In this study, seven participants with disabilities provided remote customer service at a café near the University of Tokyo using a combination of robots and personalized avatars. The results demonstrated that the avatars enabled participants to shift their identities during and after customer interactions. The study utilized longitudinal semi-structured interviews as one method to document these identity transformations. In the paper, participant P2, diagnosed with somatoform disorder, constrasted her experiences when represented by a robot and an alpaca avatar. While using the robot, P2 felt obligated to main professional behavior, adhering to the belief that "failure is not an option." Conversely, when represented by the alpaca avatar, she experienced greater freedom, allowing her to adopt more playful behavior and consequently becoming a favorite among customers. P2 noted, "With the alpaca, I can tell myself, ’It cannot be helped because I am just an alpaca.’ I cannot do that with OriHime (the robot)." She added, "When a customer says to my avatar, ’P2, go for it!’ I enjoy it because it makes me feel like I am accomplishing what I should do as an alpaca. Just walking on two legs makes me feel proud." P2 relished interacting with others as an alpaca and greatly enjoyed her identity transformation. Participant P3, who has cerebral palsy and uses a wheelchair, identifies himself as biologically female but has a fluid gender identity (male or X). P3 created an anime-style character perceived as male or gender-neutral, aligning with his gender preferences and pronouns. Typically subjected to perceptions of being "small and cute" due to his high-pitched voice and physical appearance, P3 noted significant shifts in how others treated him when using a taller, masculine avatar. This shift in perception facilitated a transformation in his self-expression, moving from the neutral Japanese first-person pronoun "watashi" to the more masculine informal "ore," and adopting a less direct but more assertive communication style. P3 described this as "finally becoming a boy," highlighting the profound impact of the avatar on his identity perception.
@@
The Avatar Robot Cafe The Avatar Robot Cafe case offers valuable insights into the intricate nature of metaverse identities. In this scenario, avatars serves as powerful tools that allow individuals to reshape their self-identity, transcend physical limitations, and engage more fully in social activities. For individual with disabilities in this case, avatars facilitate a novel identity experience and new modes of social interaction. As the metaverse continues to evolve, we believe that more such scenarios will emerge, allowing a broader and more diverse group of people to have similar experiences. These scenarios represented by this case leads us to propose the second core principle of Metaverse Identity: Fusion and Expansiveness. In the metaverse, Fusion refers to the profound integration of metaverse and real-world identities, creating an intertwined and mutually constructive holistic self. The metaverse identity anticulates needs, desires, and values that may be unexpressed or inexpressible in the real world, thus acting as an extension and expansion of the real-world identity. Conversely, the real-world identity provides foundational support for the metaverse identity, with both complementing each other. Expansiveness underscores the unprecedented possibilities for diverse expression and the infinite extension of identity in the metaverse. Individuals can transcend their inherent traits and conditions, crafting rich and varied alternative identities to present themselves in different forms and fulfill diverse needs, such as social interaction and self-realization. As illustrated by participants P2 and P3 in Avatar Robot Cafe case, who manifested new selves through an alpaca and a masculinized anime character respectively, the openness and inclusivity of the metaverse allow marginalized groups to break free from identity confinements and gain equal participation opportunities. The Fusion and Expansiveness principle further elucidates the complexity and significant of constructing identities in the metaverse. Metaverse identities are not merely collections of digital representation, data, and IDs (WEF,, 2024), but are authentic reflections of an individual’s cognition, awareness, and socio-cultural interactions. They complement and shape the subjectivity of individuals in the metaverse-era alongside real-world identities. Furthermore, they underscore the emancipatory potential of the metaverse for individuals, emphasizing how the openness, plasticity, and fluidity of identity, empowered by technology, not only offer boundless possibilities for higher-dimensional self-realization but also hold revolutionary potential for promoting social equity and fostering a diverse, inclusive ecosystem.
@@
It is noteworthy to recognize the principles of Equivalence and Alignment and Fusion and Expansiveness are complementary rather than contradictory. Equivalence and Alignment emphasizes the congruence and parity between metaverse and physical-world identities concerning behavioral norms and rights protection. This principle acts as a normative foundation, ensuring that metaverse identities do not become havens for evading responsibilities or violating others’ rights, thereby providing an essential ethical and legal framework for the integrating the metaverse and physical-world identities. Fusion and Expansiveness, on the other hand, highlights the dynamic interplay and mutual evolution between metaverse and physical-world identities through their interaction. It emphasizes the unique role of metaverse identities in shaping self-conception and pursuing values, highlighting their potential for creativity and agency. This principle serves as a catalyst for the multifaceted development of individuals and the advancement of social inclusivity. Therefore, in the governance and development of metaverse identities, it is imperative to uphold Equivalence and Alignment to regluate and guide behaviors in the metaverse, while also leveraging the potential of Fusion and Expansiveness to allow extensive exploration of identity and value creation. This dual approach necessitates the establishment of clear boundaries and behavioral guidelines for metaverse and real-world identities when formulating pertinent policies and regulations, while preserving sufficient flexibility and openness to encourage the expression and integration of diverse identities, thereby fostering social inclusively and innovation. Only through the dynamic equilibrium of these two principles can we realize a synergistic interaction between digital and real coexistence, ensuring the orderly operation of the metaverse ecosystem while unlocking the positive potential of technological advancement.
@@ 4.1 Identity Interoperability Challenges
Interoperability challenges are a critical obstacle to the development of the metaverse, as discussed by both academia and industry in recent years (Ball,, 2022; Wang et al.,, 2022; WEF, 2023a, ). In the systematic literature review on metaverse interoperability, Yang et al., 2024a emphasize that identity is the cornerstone of an interconnected Metaverse. Without a consistent and seamless identity transfer between the physical world and the metaverse, as well as across different platforms within the metaverse, true inter-connectivity remains unattainable. Equivalence and Alignment stress the consistency and equal status of metaverse identities and real-world identities in the terms of behavioral norms and rights protections. Just as our identities in the physical world an seamlessly transition, authenticate, and connect across different scenarios, metaverse identities should exhibit the same fluidity. This is essential for the metaverse to be a true replica and extention of the physical world (Wang et al.,, 2022).
@@
Following the three components framework (WEF,, 2024), the current interoperable state of metaverse identity is as follows: At the identity representation layer, most metaverse platforms continue the practices of the internet era, requiring users to create separate avatars on different platforms (Jiang et al.,, 2023). At the data layer, users generate identity-related data such as personal attributes, social relationships, and behavioral records on various platforms, but this data is typically stored and managed in isolated silos by each platform (Wang et al.,, 2022), preventing effective cross-domain data management and sharing. Finally, at the identification layer, centralized methods dominate despite exploratory applications such decentralized identity (DID) technologies, which remain immature stages with significant shortages (Alizadeh et al.,, 2022). This results in limited interoperability due to varying DID solutions across platforms, and the coexistence of centralized and decentralized methods complicates interconnected identification (Yang et al., 2024a, ), necessitating further resolution. These fragmented identities lack the unified digital identity mappings or authorization mechanisms, necessitating repeated customization and offering minimal support for mutual recognition and migration across platforms, which disrupts user identity consistency.
@@
Although the nascent stage of the metaverse’s development is a primary factor contributing to these issues, it is imperative to consider the risks associated with path dependency in advance. The prevalent identity systems are predominantly centralized and designed around individual platforms, inherently lacking interoperability. Economic incentives often drive platforms to implement "walled garden" strategies, thus inhibiting the advancement of interoperability (Kerber and Schweitzer,, 2017). Furthermore, the absence of mature technical standards and subsequent legal frameworks constrains progress in this domain. It is also essential to acknowledge that the principle of Fusion and Expansiveness underscore the diversity and dynamic nature of metaverse identities, which can lead to more complex scenarios due to interoperability challenges.
@@ 4.2 Legal Challenges of Identity in Metaverse
As the metaverse continues to evolve and its applications expand, the challenges to legal boundaries become increasingly evident. In the cases of sexual assault discussed previously, the victims experienced substantial psychological harassment. Current legal frameworks, however, find it difficult to enforce traditional accountability and sanctions due to the lack of physical contact (Cheong,, 2022; Lee et al., 2023b, ; Bellini,, 2024). The report, "Sexual Violence and Harassment in the Metaverse: A New Manifestations of Gender-Based Harms," underscores the urgent need for updated national and international law enforcement to address this distinct form of metaverse-facilitated sexual violence and harassment (Rigotti and Malgieri,, 2024). With avatar-based interactions becoming commonplace in the metaverse, the frequency of such rights conflicts and illegal acts is expected to rise. Actions that would typically incur legal penalties in the real world, from civil liabilities like negligence and harassment to criminal offenses such as assault and theft, pose significant challenges in virtual settings of metaverse. The applicability and efficacy of existing laws within these virtual contexts are subjects of ongoing significant debate (Cheong,, 2022; Lee et al., 2023b, ).
@@
Existing legal frameworks struggle to address issues of identity recognition, behavior evaluation, and accountability when these entities engage in detrimental activities (WEF,, 2024). One proposed solution from the academic realm to address the challenges posed by digital entities is to grant them the status of legal personhood, thus holding them accountable for their actions within the metaverse (Kurki,, 2019). This would involve incorporating rights and responsibilities for these entities within the existing legal framework and granting them the capacity to be party to legal proceedings, which introduces significant complexity. Key issues include defining legal standards for the recognition of digital entities and managing the effects of their actions on both their controllers and other stakeholders, necessitating innovative legal and policy responses (Cheong,, 2022).

### v2

@@ 3.1 Equivalence and Alignment
In recent years, the incidence of sexual assaults within the metaverse has become as a pressing concern, presenting complex social and legal dilemmas (Schulenberg et al.,, 2023; Lee et al., 2023b, ; Bellini,, 2024). For instance, in December 2021, a female Metaverse user reported that three male avatars had ‘touched her[her avatar] inappropriately’ (Clayton,, 2022). Another disturbing incident occurred in April 2022, when a female user from Japan shared via Twitter her experience of being attacked and virtually raped in VRChat while she was asleep (Anonymous,, 2022). Additionally, the Daily Mail in the UK highlighted a particularly alarming case in January with its headline “British police probe Virtual Rape in metaverse: Young girl’s digital persona ‘is sexually attacked by gang of adult men in immersive video game’ - sparking the first investigation of its kind and questions about extent current laws apply in online world” (Camber,Rebecca,, 2024). The incident involved a British girl under the age of 16 whose avatar was raped by multiple avatars controlled by strangers during a virtual-reality game, leading to the first global investigation of such behaviors and raising significant legal inquiries regarding the applicability of existing laws in the metaverse.
@@
At first glance, these cases may seem to involve merely disputes over unethical conduct and criminal actions, along with attendant legal applicability issues. However, a deeper analysis reveals that they fundamentally reflect profound challenges associated with identity recognition in the metaverse (Cheong,, 2022; Kasiyanto et al.,, 2022; Wu and Zhang,, 2023; Lee et al., 2023b, ). The reason these incidents of sexual assault attract world-spread social attention and provoke strong reaction lies in the deep psychological identification that individuals have with their avatars (Lee et al., 2023b, ). If avatars were merely considered gaming tools or entertainment elements, violations against them would not significantly impact users in reality. Nevertheless, the victims’ report of discomfort and psychological trauma suggest a profound association between their avatar and their identities. This strong identification means that harm inflicted on an avatar is perceived as harm to the self, eliciting genuine emotional and psychological responses (Lee et al., 2023b, ). The legal disputes triggered by these cases also underscore the unique nature of metaverse identities in both legal and social contexts. While violations of personal rights in the physical world inevitably entail legal repercussions, in the metaverse, the identity attributes of avatars remain undefined, leading to ambiguities in rights protection (Lee et al., 2023b, ; Bellini,, 2024). This ambiguity necessitates a redefinition of the identities represented by avatars. Such incidents of ‘sexual assault’ not only reveal the high degree of integration between avatars and users’ self-presence-a psychological state in which people equate their virtual selves with their actual selves (Lee et al., 2023b, )- but also demonstrate the significant influence of metaverse identities on real-life identity, highlighting the close connection between the metaverse and real life.
@@
It is critical to acknowledge that the identity dilemmas triggered by these ‘sexual assault’ cases represent just one of the numerous challenges posed by metaverse entities in the metaverse. OpenAI’s recent release of ChatGPT-4o [^2] https://openai.com/index/hello-gpt-4o/, an AI chatbot capable of human-like interaction, introduces a novel and complex dimension of challenges for identity recognition in the metaverse (Hsu,Jeremy,, 2024). In this context, we propose the core principle - Equivalence and Alignment for conceptualizing and governing Metaverse Identity. Alignment refers to the consistency between metaverse identities and real-world identities concerning behavioral norms, social guidelines, and ethical standards. Equivalence emphasizes the parity between the two in aspects such as behavioral patterns, emotional experiences, cognitive perceptions, and economic interests (Lee et al., 2023b, ). Essentially, Alignment mandates that metaverse identities adheres to the same standards as real-world identities, while Equivalence underscores that the impact of virtual identities on users is as significant as that of real-world identities. We aim to simplify the complex attribute of Metaverse Identity through this principle, providing a clear and straightforward framework for developing metaverse identities. The principle also underscores the unique status of Metaverse Identity in contemporary legal and social contexts in metaverse era compared to in internet era, indicating the necessity of refining relevant regulations in the physical world. Furthermore, we intend to offer principled recommendations for policymakers, technology developers, and platform operators. While striving for deeper immersion and realism in the metaverse, it is imperative to prioritize the recognition and protection of user identities, thereby guiding necessary regulation and ethical oversight at the intersection of technology and ethics.
@@ 3.2 Fusion and Expansiveness
Hatada et al., (2024) recently published their study “Avatar Robot Cafe,” which explores how individuals with disabilities employ robots, avatars and a hybrid cyber-physical environment to redefine their identities. In this study, seven participants with disabilities provided remote customer service at a café near the University of Tokyo using a combination of robots and personalized avatars. The results demonstrated that the avatars enabled participants to shift their identities during and after customer interactions. The study utilized longitudinal semi-structured interviews as one method to document these identity transformations. In this work, participant P2, diagnosed with somatoform disorder, constrasted her experiences when represented by a robot and an alpaca avatar. While using the robot, P2 felt obligated to main professional behavior, adhering to the belief that “failure is not an option.” Conversely, when represented by the alpaca avatar, she experienced greater freedom, allowing her to adopt more playful behavior and consequently becoming a favorite among customers. P2 noted, “With the alpaca, I can tell myself, ‘It cannot be helped because I am just an alpaca.’ I cannot do that with OriHime (the robot).” She added, “When a customer says to my avatar, ‘P2, go for it!’ I enjoy it because it makes me feel like I am accomplishing what I should do as an alpaca. Just walking on two legs makes me feel proud.” P2 relished interacting with others as an alpaca and greatly enjoyed her identity transformation. Participant P3, who has cerebral palsy and uses a wheelchair, identifies himself as biologically female but has a fluid gender identity (male or X). P3 created an anime-style character perceived as male or gender-neutral, aligning with his gender preferences and pronouns. Typically subjected to perceptions of being “small and cute” due to his high-pitched voice and physical appearance, P3 noted significant shifts in how others treated him when using a taller, masculine avatar. This shift in perception facilitated a transformation in his self-expression, moving from the neutral Japanese first-person pronoun “watashi” to the more masculine informal “ore,” and adopting a less direct but more assertive communication style. P3 described this as “finally becoming a boy,” highlighting the profound impact of the avatar on his identity perception.
@@
The Avatar Robot Cafe case offers valuable insights into the intricate nature of metaverse identities. In this scenario, avatars serves as powerful tools that allow individuals to reshape their self-identity, transcend physical limitations, and engage more fully in social activities. For individual with disabilities in this case, avatars facilitate a novel identity experience and new modes of social interaction. As the metaverse continues to evolve, we believe that more such scenarios will emerge, allowing a broader and more diverse group of people to have similar experiences. These scenarios represented by this case leads us to propose the second core principle of Metaverse Identity: Fusion and Expansiveness. Within this context, Fusion refers to the profound integration of metaverse and real-world identities, creating an intertwined and mutually constructive holistic self. The metaverse identity articulates needs, desires, and values that may be unexpressed or inexpressible in the real world, thus acting as an extension and expansion of the real-world identity. Conversely, the real-world identity provides foundational support for the metaverse identity, with both complementing each other. Expansiveness underscores the unprecedented possibilities for diverse expression and the infinite extension of identity in the metaverse. Individuals can transcend their inherent traits and conditions, crafting rich and varied alternative identities to present themselves in different forms and fulfill diverse needs, such as social interaction and self-realization.
As illustrated by participants P2 and P3 in Avatar Robot Cafe case, who manifested new selves through an alpaca and a masculinized anime character respectively, the openness and inclusivity of the metaverse allow marginalized groups to break free from identity confinements and gain equal participation opportunities. The Fusion and Expansiveness principle further elucidates the complexity and significant of constructing identities in the metaverse. Metaverse identities are not merely collections of digital representation, data, and IDs (WEF,, 2024), but are authentic reflections of an individual’s cognition, awareness, and socio-cultural interactions. They complement and shape the subjectivity of individuals in the metaverse-era alongside real-world identities. Furthermore, they underscore the emancipatory potential of the metaverse for individuals, emphasizing how the openness, plasticity, and fluidity of identity, empowered by technology, not only offer boundless possibilities for higher-dimensional self-realization but also hold revolutionary potential for promoting social equity and fostering a diverse, inclusive ecosystem.
@@
It is noteworthy to recognize the principles of Equivalence and Alignment and Fusion and Expansiveness are complementary rather than contradictory. Equivalence and Alignment emphasizes the congruence and parity between metaverse and physical-world identities concerning behavioral norms and rights protection. This principle acts as a normative foundation, ensuring that metaverse identities do not become havens for evading responsibilities or violating others’ rights, thereby providing an essential ethical and legal framework for the integrating the metaverse and physical-world identities. Fusion and Expansiveness, on the other hand, highlights the dynamic interplay and mutual evolution between metaverse and physical-world identities through their interaction. It emphasizes the unique role of metaverse identities in shaping self-conception and pursuing values, highlighting their potential for creativity and agency. This principle serves as a catalyst for the multifaceted development of individuals and the advancement of social inclusivity. Therefore, in the governance and development of metaverse identities, it is imperative to uphold Equivalence and Alignment to regulate and guide behaviors in the metaverse, while also leveraging the potential of Fusion and Expansiveness to allow extensive exploration of identity and value creation. This dual approach necessitates the establishment of clear boundaries and behavioral guidelines for metaverse and real-world identities when formulating pertinent policies and regulations, while preserving sufficient flexibility and openness to encourage the expression and integration of diverse identities, thereby fostering social inclusively and innovation. Only through the dynamic equilibrium of these two principles can we realize a synergistic interaction between digital and real coexistence, ensuring the orderly operation of the metaverse ecosystem while unlocking the positive potential of technological advancement.
@@ 4.1 Identity Interoperability Challenges
Interoperability challenges are a critical obstacle to the development of the metaverse, as discussed by both academia and industry in recent years (Ball,, 2022; Wang et al.,, 2022; WEF, 2023a, ). In the systematic literature review on metaverse interoperability, Yang et al., 2024a emphasize that identity is the cornerstone of an interconnected Metaverse. Without a consistent and seamless identity transfer between the physical world and the metaverse, as well as across different platforms within the metaverse, true inter-connectivity remains unattainable. Equivalence and Alignment stress the consistency and equal status of metaverse identities and real-world identities in the terms of behavioral norms and rights protections. Just as our identities in the physical world a seamlessly transition, authenticate, and connect across different scenarios, metaverse identities should exhibit the same fluidity. This is essential for the metaverse to be a true replica and extention of the physical world (Wang et al.,, 2022).
@@
Following the three components framework (WEF,, 2024), the current interoperable state of metaverse identity is as follows: At the identity representation layer, most metaverse platforms continue the practices of the internet era, requiring users to create separate avatars on different platforms (Jiang et al.,, 2023). At the data layer, users generate identity-related data such as personal attributes, social relationships, and behavioral records on various platforms, but this data is typically stored and managed in isolated silos by each platform (Wang et al.,, 2022), preventing effective cross-domain data management and sharing. Finally, at the identification layer, centralized methods dominate despite exploratory applications such as decentralized identity (DID) technologies, which remain in immature stages with significant shortages (Alizadeh et al.,, 2022). This results in limited interoperability due to varying DID solutions across platforms, and the coexistence of centralized and decentralized methods complicates interconnected identification (Yang et al., 2024a, ), necessitating further resolution. These fragmented identities lack unified digital identity mappings or authorization mechanisms, necessitating repeated customization and offering minimal support for mutual recognition and migration across platforms, which disrupts user identity consistency.
@@
Although the nascent stage of the metaverse’s development is a primary factor contributing to these issues, it is imperative to consider the risks associated with path dependency in advance. The prevalent identity systems are predominantly centralized and designed around individual platforms, inherently lacking interoperability. Economic incentives often drive platforms to implement “walled garden” strategies, thus inhibiting the advancement of interoperability (Kerber and Schweitzer,, 2017). Furthermore, the absence of mature technical standards and subsequent legal frameworks constrains progress in this domain (Lee et al., 2023b, ; Yang et al., 2024a, ). It is also essential to acknowledge that the principle of Fusion and Expansiveness underscore the diversity and dynamic nature of metaverse identities, which can lead to more complex scenarios due to interoperability challenges.
@@ 4.2 Legal Challenges of Identity in Metaverse
As the metaverse continues to evolve and its applications expand, the challenges to legal boundaries become increasingly evident. In the cases of sexual assault discussed previously, the victims experienced substantial psychological harassment. Current legal frameworks, however, find it difficult to enforce traditional accountability and sanctions due to the lack of physical contact (Cheong,, 2022; Lee et al., 2023b, ; Bellini,, 2024). The report, “Sexual Violence and Harassment in the Metaverse: A New Manifestations of Gender-Based Harms,” underscores the urgent need for updated national and international law enforcement to address this distinct form of metaverse-facilitated sexual violence and harassment (Rigotti and Malgieri,, 2024). With avatar-based interactions becoming commonplace in the metaverse, the frequency of such rights conflicts and illegal acts is expected to rise. Actions that would typically incur legal penalties in the real world, from civil liabilities like negligence and harassment to criminal offenses such as assault and theft, pose significant challenges in virtual settings of metaverse. The applicability and efficacy of existing laws within these virtual contexts are subjects of ongoing significant debate (Cheong,, 2022; Lee et al., 2023b, ).
@@
Existing legal frameworks struggle to address issues of identity recognition, behavior evaluation, and accountability when these entities engage in detrimental activities (WEF,, 2024). One proposed solution from the academic realm to address the challenges posed by digital entities is to grant them the status of legal personhood, thus holding them accountable for their actions within the metaverse (Kurki,, 2019). This would involve incorporating rights and responsibilities for these entities within the existing legal framework and granting them the capacity to be a party to legal proceedings, which introduces significant complexity. Key issues include defining legal standards for the recognition of digital entities and managing the effects of their actions on both their controllers and other stakeholders, necessitating innovative legal and policy responses (Cheong,, 2022).